OpenStudy (mlgpro):

Which set of line segments with the given lengths can be formed into a right triangle? twenty four, thirty, thirty five twelve, eighteen, thirty eighteen, twenty four, thirty eighteen, twenty four, thirty five

OpenStudy (anonymous):

use pythagorean theoreom in all of them

OpenStudy (anonymous):

lol that will be a long way but you will get the answer.

OpenStudy (mlgpro):

i cant

OpenStudy (anonymous):

why?

OpenStudy (mlgpro):

i dont know how

OpenStudy (anonymous):

18^2+24^2=30^2 324+576=900 900=900 :)

OpenStudy (anonymous):

do you understand what I did?

OpenStudy (anonymous):

I plug in those values in the formula of pythagorean theorem

OpenStudy (mlgpro):

yes you kinda gave me the answer

OpenStudy (anonymous):

well tell me what you don't understand about it?.

OpenStudy (anonymous):

The only thing I did was let a=the first value, b= the second value,and c=third value

OpenStudy (mlgpro):

my mom rild me it was just a^2+b^2=c^2 and i only got 1 number

OpenStudy (anonymous):

what do you mean?

OpenStudy (mlgpro):

i only got c^2

OpenStudy (anonymous):

you're suppose to get only one number in the end a^2 and B^2 will add together for B^2

OpenStudy (anonymous):

C^2*

OpenStudy (anonymous):

yes!. they have to be equal

OpenStudy (anonymous):

A will always be the first term given then B will be the second term given and C will be the second term given unless they give A= C= now find B then you would square both of them and divide A into C to find B

OpenStudy (anonymous):

C will be the third* can't type today

OpenStudy (mlgpro):

ok